# Differentially Private Clustered Federated Learning

### v5

@@ C.4 Privacy Budgeting
In order to have a fair comparison between our algorithm and the baselines, we align them all to have the same “total” privacy budget $\epsilon$ and satisfy $(\epsilon,\delta)$-DP for a fixed $\delta$. In order to account the privacy of an algorithm, we compose the RDP guarantees of all private operations in the algorithm and then convert the resulting RDP guarantee to approximate $(\epsilon,\delta)$-DP using C.2. The DPSGD performed by different algorithms for local training benefits from privacy amplification by subsampling (Mironov et al., 2019). Algorithms that have privacy overheads, e.g., IFCA and R-DPCFL which need to privatize their local clustering as well, will have less privacy budget left for training. In other words, for the same total privacy budget $\epsilon$, IFCA and R-DPCFL will use a larger amount of noise when running DPSGD, compared to MR-MTL that has zero privacy overhead.
@@ D.6 Implementation of private local clustering for IFCA and R-DPCFL
In every round of IFCA and during the rounds $e>E_{c}$ of R-DPCFL, the server sends $M$ cluster models to all clients, and they evaluate them on their local datasets. Then, each client $i$ selects the model with the lowest loss on its local dataset $\mathcal{D}_{i}$, trains it for $K$ local epochs and sends the result back to the server. This model selection performed by each client can lead to privacy leakage w.r.t its local dataset, if it is not privatized. In order to protect data privacy, clients need to privatize their local clustering by using exponential mechanism and accounting its privacy using z-CDP, explained in Section C.3. Assuming a total privacy budget $\epsilon$ for a client $i$, it has to split the budget between private clustering and DPSGD. Naive split of privacy budget can lead to very noisy DPSGD steps or very noisy local selection by clients. Following (Liu et al., 2022a), we use two strategies to mitigate the privacy overhead of local clustering performed by IFCA and R-DPCFL:
@@ D.8 Gaussian Mixture Model
1) parameter initialization, which we set to “k-means++”. This is because this type of initialization leads to both low time to initialize and low number of EM iterations for the GMM to converge
@@ E.3 The number of clusters $M$
Appendix F Complete experimental results
The following figures, which include the results for the Global baseline, are the more complete versions of the figures in the paper (Figure 4 and Figure 5).
@@ Appendix F Complete experimental results
The following figure shows how the MSS score of the learned GMM model at the end of the first round can be indicative of whether the true clients’ clusters will be detected or not. As observed, an MSS score above 2 almost always yields to correct detection of all clusters.
@@ G.4 Formal privacy guarantees of R-DPCFL
The privacy guarantee of R-DPCFL for each client $i$ in the system comes from the fact that the client runs DPSGD with a fixed DP noise variance $\sigma_{i,\texttt{DP}}^{2}=c^{2}\cdot z_{i}^{2}(\epsilon,\delta,b_{i}^{1},b_{i}^{>1},N_{i},K,E)$ in each of its batch gradient computations. We provide a formal privacy guarantee for the algorithm to show the sample-level DP privacy guarantees provided to each client $i$ with respect to its local dataset $\mathcal{D}_{i}$ and against the untrusted server (and any other external third party).
@@
The sensitivity of the batch gradient in Equation 1 to every data sample is $c$. Therefore, based on Proposition 7 in (Mironov, 2017) each of the batch gradient computations by client $i$ (in the first round $e=1$ as well as the next rounds $e>1$) is $(\alpha,\frac{\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}})$-RDP w.r.t the local dataset $\mathcal{D}_{i}$. Therefore, if the client runs $E_{i}^{\texttt{ tot}}$ total number of gradient updates during the training time, which results in the model updates $\{\Delta\tilde{\bm{\theta}}_{i}^{e}\}_{e=1}^{E}$ uploaded to the server, the set of model updates will be $(\alpha,\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}})$-RDP w.r.t $\mathcal{D}_{i}$, according to Proposition 1 in (Mironov, 2017). Finally, according to Proposition 3 in the same work, this guarantee is equivalent to $(\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}}+\frac{log(1/\delta)}{\alpha-1},\delta)$-DP (for any $\delta>1$). The RDP-based guarantee is computed over a bunch of orders $\alpha$ and the best result among them is chosen as the privacy guarantee. Therefore, the proof is complete and the set $\{\Delta\tilde{\bm{\theta}}_{i}^{e}\}_{e=1}^{E}$ satisfies $(\epsilon,\delta)$-DP w.r.t $\mathcal{D}_{i}$, with $\epsilon=\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}}+\frac{log(1/\delta)}{\alpha-1}$ derived above, and $\delta>0$. Tight bounds for $\epsilon$ can be derived by using the numerical procedure, proposed in (Mironov et al., 2019), for accounting sampled Gaussian mechanism. On the other hand, clients’ local cluster selections are also privatized by exponential mechanism and satisfy $(\epsilon,\delta)$-DP. Therefore, the overall training process for each client is private and satisfies ($\epsilon,\delta$)-DP. ∎

### v6

@@ C.4 Privacy Budgeting
In order to have a fair comparison between our algorithm and the baselines, we align them all to have the same “total” privacy budget $\epsilon$ and satisfy $(\epsilon,\delta)$-DP for a fixed $\delta$. In order to account the privacy of an algorithm, we compose the RDP guarantees of all private operations in the algorithm and then convert the resulting RDP guarantee to approximate $(\epsilon,\delta)$-DP using C.3. The DPSGD performed by different algorithms for local training benefits from privacy amplification by subsampling (Mironov et al., 2019). Algorithms that have privacy overheads, e.g., IFCA and R-DPCFL which need to privatize their local clustering as well, will have less privacy budget left for training. In other words, for the same total privacy budget $\epsilon$, IFCA and R-DPCFL will use a larger amount of noise when running DPSGD, compared to MR-MTL that has zero privacy overhead.
@@ D.6 Implementation of private local clustering for IFCA and R-DPCFL
In every round of IFCA and during the rounds $e>E_{c}$ of R-DPCFL, the server sends $M$ cluster models to all clients, and they evaluate them on their local datasets. Then, each client $i$ selects the model with the lowest loss on its local dataset $\mathcal{D}_{i}$, trains it for $K$ local epochs and sends the result back to the server. This model selection performed by each client can lead to privacy leakage w.r.t its local dataset, if it is not privatized. In order to protect data privacy, clients need to privatize their local clustering by using exponential mechanism and accounting its privacy using z-CDP, explained in Section C.3. Assuming a total privacy budget $\epsilon$ for a client $i$, it has to split the budget between private clustering and DPSGD. Naive split of privacy budget can lead to very noisy DPSGD steps or very noisy local cluster selection by clients. Following (Liu et al., 2022a), we use two strategies to mitigate the privacy overhead of local clustering performed by IFCA and R-DPCFL:
@@ D.8 Gaussian Mixture Model
1) parameter initialization, which we set to “k-means++”. This is because this type of initialization leads to both low time to initialize and low number of EM iterations for the GMM to converge (Arthur & Vassilvitskii, 2007; Biernacki et al., 2003).
@@ E.3 The number of clusters $M$
Appendix F More experimental results
The results shown in Figure 10 and Figure 11 include the results for the Global baseline and are the more complete versions of the figures in the paper (Figure 4 and Figure 5).
@@ Appendix F More experimental results
Figure 12 shows how the MSS score of the learned GMM at the first round can be indicative of whether the true clients’ clusters will be detected correctly or not. An MSS score above 2 almost always yields to correct detection of all clusters.
@@
The results in Figure 13 supports Theorem 4.3 by showing that if $b_{i}^{1}$ is large enough, the convergence rate of EM algorithm for learning the GMM at the end of the first round increases with $b_{i}^{1}$.
Figure 13: The number of EM iterations needed for learning the GMM decreases as $b_{i}^{1}$ increases. Especially, at full batch size $b_{i}^{1}=N_{i}(\forall i)$, very few iterations of EM are needed. The results are obtained on CIFAR10 with covariate shift across $M=4$ clusters. According to 4.2, using a large enough batch size $b_{i}^{1}$ in the first round makes the underlying clusters in $\{\Delta\tilde{\bm{\theta}}_{i}^{1}\}_{i=1}^{n}$ more distinguishable. Consequently, according to Theorem 4.3, convergence rate of EM algorithm for learning the GMM increases with $b_{i}^{1}$.
@@ G.4 Formal privacy guarantees of R-DPCFL
The privacy guarantee of R-DPCFL for each client $i$ in the system comes from the fact that the client runs DPSGD with a fixed DP noise variance $\sigma_{i,\texttt{DP}}^{2}=c^{2}\cdot z_{i}^{2}(\epsilon,\delta,b_{i}^{1},b_{i}^{>1},N_{i},K,E)$ in each of its batch gradient computations. In the following Theorem G.1, we provide a formal privacy guarantee for the algorithm to show the sample-level DP privacy guarantees provided to each client $i$ with respect to its local dataset $\mathcal{D}_{i}$ and against the untrusted server (and any other external third party).
@@
The sensitivity of the batch gradient in Equation 1 to every data sample is $c$. Therefore, based on Proposition C.4, each of the batch gradient computations by client $i$ (in the first round $e=1$ as well as the next rounds $e>1$) is $(\alpha,\frac{\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}})$-RDP w.r.t the local dataset $\mathcal{D}_{i}$. Therefore, if the client runs $E_{i}^{\texttt{ tot}}$ total number of gradient updates during the training time, which results in the model updates $\{\Delta\tilde{\bm{\theta}}_{i}^{e}\}_{e=1}^{E}$ uploaded to the server, the set of model updates will be $(\alpha,\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}})$-RDP w.r.t $\mathcal{D}_{i}$, according to Proposition C.2. Finally, according to C.3, this guarantee is equivalent to $(\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}}+\frac{log(1/\delta)}{\alpha-1},\delta)$-DP (for any $\delta>1$). The RDP-based guarantee is computed over a bunch of orders $\alpha$ and the best result among them is chosen as the privacy guarantee. Therefore, the proof is complete and the set $\{\Delta\tilde{\bm{\theta}}_{i}^{e}\}_{e=1}^{E}$ satisfies $(\epsilon,\delta)$-DP w.r.t $\mathcal{D}_{i}$, with $\epsilon=\frac{E_{i}^{\texttt{ tot}}\alpha c^{2}}{2\sigma_{i,\texttt{DP}}^{2}}+\frac{log(1/\delta)}{\alpha-1}$ derived above, and $\delta>0$. Tight bounds for $\epsilon$ can be derived by using the numerical procedure, proposed in (Mironov et al., 2019), for accounting sampled Gaussian mechanism. On the other hand, clients’ local cluster selections are also privatized by exponential mechanism and satisfy $(\epsilon,\delta)$-DP. Therefore, the overall training process for each client is private and satisfies ($\epsilon,\delta$)-DP. ∎